OpenStudy (anonymous):

Write the equation of the quadratic function with roots -2 and 4 and a vertex at (1, 9).

OpenStudy (anonymous):

y=(x+2)(x-4) When x=1, y=9 Sub into the equation x=1 and y=9 \[(1+2)(1-4)=9\] \[(3)(-3)=9\] \[-9=9\] The LHS(Left Hand Side) must equal the RHS(Right Hand Side). That means you have to multiply -9 by -1. Therefore the equation would have a minus sign in front of the equation. \[y=-(x+2)(x-4)\]